OpenStudy (anonymous):

Find the area bounded by a^2y=x^3,the x-axis and the line x=2a

OpenStudy (jack1):

umm... is integral area under the line...? i honestly forget, standby... @terenzreignz

OpenStudy (anonymous):

its a topic related to uses of antiderivatives

OpenStudy (anonymous):

its above the line

OpenStudy (anonymous):

\[Area =\int\limits_{a}^{b}(y)dx\]

OpenStudy (anonymous):

this is the only thing i know

OpenStudy (anonymous):

i can't sketch it . i dont even know which area it is referring to

OpenStudy (jack1):

x = 2a so a = x/2 so equation is (x/2)^2y = x^3 ...yeah?

OpenStudy (anonymous):

yeah

terenzreignz (terenzreignz):

a is just a constant, right? :D

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

the answer it shows at the back of my textbook is 4a^2 units^2

terenzreignz (terenzreignz):

So... this basically ensures that a^2 is positive. Well then, let's sketch that bloody graph XD \[\Large y = \frac{x^3}{a^2}\]